# Real-World Remote Sensing Image Dehazing: Benchmark and Baseline

### v1

@@ V-C3 Capacity of CSAM
TABLE VII: Performance comparison of different fusion modules
Component PSNR SSIM MSE #Param FLOPs
$1\times 1$ Conv 25.01 0.7128 0.0045 563.1k 20.02G
SKFusion [7] 25.19 0.7132 0.0043 548.5k 19.42G
MFAFM 25.40 0.7221 0.0040 558.1K 19.82G
@@
TABLE VIII: Performance comparison of different fusion modules
@@ V-C4 Adaptive Feature Fusion
To comprehensively validate the efficacy of MFAFM, we performed systematic comparisons with two representative fusion approaches: standard 1$\times$1 convolution and SKFusion [7]. As quantitatively demonstrated in Table VIII, the proposed MFAFM achieves marked superiority over conventional convolution operations, achieving a significant gain in PSNR of 0.39 dB. Although SKFusion marginally outperforms baseline convolution (0.12 dB improvement), our method established a new SOTA by further increasing the performance ceiling with an additional 0.21 dB enhancement. This progression not only corroborated the findings in [7] but more importantly underscores MFAFM’s exceptional capability in addressing real-world RSID challenges, confirming MFAFM’s effectiveness as a multiscale adaptive paradigm for feature fusion in RSID applications.
@@ V-C5 Loss Functions
To investigate the effectiveness of different loss functions, we conducted ablation studies with various loss combinations. Table IX presented the quantitative results. Using only $\mathcal{L}_{2}$ loss achieved a PSNR of 24.95 dB. When replacing it with perpetual loss, the PSNR decreased to 22.85 dB, but showed better perceptual quality with a SSIM of 0.6873. By combining $\mathcal{L}_{2}$ and $\mathcal{L}_{perpetual}$ losses, our model achieved the best performance with a PSNR of 25.40 dB and SSIM of 0.7221, demonstrating the complementary effects of these two loss terms in optimizing both pixel-level accuracy and perceptual quality.
TABLE IX: Performance comparison of different loss functions.
@@
$\mathcal{L}_{perpetual}$ 22.85 0.6873 0.0052
$\mathcal{L}_{2}+\lambda\mathcal{L}_{perpetual}$ 25.40 0.7221 0.0040
@@ V-D Efficiency Analysis
We conducted a comprehensive comparison of SOTA dehazing methods in Table X, evaluating their restoration quality and computational efficiency. MCAF-Net established new benchmarks for RSID by simultaneously achieving SOTA restoration quality (PSNR 25.40 dB, SSIM 0.7221) and unparalleled efficiency (558.1K parameters, 19.82G FLOPs). While Trinity-Net processes images marginally faster (0.0400s vs. 0.0498s), its PSNR lags by 1.29 dB. Compared to DehazeFormer, the prior quality leader, our method reduces the inference time by 43.9% (0.0831s vs. 0.0498s) while improving the PSNR by 0.71 dB and using 53.7% fewer parameters. This dual optimization of accuracy and resource efficiency positioned MCAF-Net as the optimal choice for real-time haze removal in resource-constrained environments.
TABLE X: Performance comparison of state-of-the-art methods
@@ VI Conclusion and Future Works
In this work, we present MCAF-Net, an innovative framework for real-world RSID, designed to overcome the limitations of synthetic-to-real domain gaps through a multi-branch adaptive architecture. Our approach integrates MFIBA, CSAM, and MFAFM to achieve comprehensive feature learning, precise color restoration, and context-aware haze suppression. The introduction of the RRSHID dataset—the first large-scale real-world benchmark—provides critical support for modeling complex atmospheric variations and sensor-specific degradations. Extensive validation demonstrates that MCAF-Net not only outperforms existing methods in real-world scenarios but also maintains competitive efficiency with minimal computational overhead, offering practical value for time-sensitive earth observation tasks. While the current framework excels in static haze removal, challenges remain in handling dynamic atmospheric interactions and multi-modal sensor data. Future efforts will focus on integrating temporal modeling and cross-spectral fusion to advance toward all-weather, physics-aware RS restoration systems.

### v2

@@ V-C2 Basic Learning Blocks
V-C3 Capacity of CSAM
To comprehensively evaluate the effectiveness of the CSAM, we conducted detailed ablation studies focusing on its two core components: the self-supervised color correction mechanism and the attention-guided feature refinement, with the results summarized in Table VIII.
@@ V-C3 Capacity of CSAM
TABLE IX: Performance comparison of different fusion modules
@@ V-C4 Adaptive Feature Fusion
To comprehensively validate the efficacy of MFAFM, we performed systematic comparisons with two representative fusion approaches: standard 1$\times$1 convolution and SKFusion [6]. As quantitatively demonstrated in Table IX, the proposed MFAFM achieves marked superiority over conventional convolution operations, achieving a significant gain in PSNR of 0.39 dB. Although SKFusion marginally outperforms baseline convolution (0.12 dB improvement), our method established a new SOTA by further increasing the performance ceiling with an additional 0.21 dB enhancement. This progression not only corroborated the findings in [6] but more importantly underscored MFAFM’s exceptional capability in addressing real-world RSID challenges, confirming MFAFM’s effectiveness as a multiscale adaptive paradigm for feature fusion in RSID applications.
@@ V-C5 Loss Functions
To investigate the effectiveness of different loss functions, we conducted ablation studies with various loss combinations. Table X presented the quantitative results. Using only $\mathcal{L}_{2}$ loss achieved a PSNR of 24.95 dB. When replacing it with perceptual loss, the PSNR decreased to 22.85 dB, but showed better perceptual quality with a SSIM of 0.6873. By combining $\mathcal{L}_{2}$ and $\mathcal{L}_{perceptual}$ losses, our model achieved the best performance with a PSNR of 25.40 dB and SSIM of 0.7221, demonstrating the complementary effects of these two loss terms in optimizing both pixel-level accuracy and perceptual quality.
TABLE X: Performance comparison of different loss functions.
@@
$\mathcal{L}_{perceptual}$ 22.85 0.6873 0.0052
$\mathcal{L}_{2}+\lambda\mathcal{L}_{perceptual}$ 25.40 0.7221 0.0040
@@ V-D Efficiency Analysis
We conducted a comprehensive comparison of SOTA dehazing methods in Table XI, evaluating their restoration quality and computational efficiency. MCAF-Net established new benchmarks for RSID by simultaneously achieving SOTA restoration quality (PSNR 25.40 dB, SSIM 0.7221) and unparalleled efficiency (558.1K parameters, 19.82G FLOPs). While Trinity-Net processes images marginally faster (0.0400s vs. 0.0498s), its PSNR lags by 1.29 dB. Compared to DehazeFormer, the prior quality leader, our method reduces the inference time by 43.9% (0.0831s vs. 0.0498s) while improving the PSNR by 0.71 dB and using 53.7% fewer parameters. This dual optimization of accuracy and resource efficiency positioned MCAF-Net as the optimal choice for real-time haze removal in resource-constrained environments.
TABLE XI: Performance comparison of state-of-the-art methods
@@ VI Conclusion and Future Works
In this work, we present MCAF-Net, an innovative framework for real-world RSID, designed to overcome the limitations of synthetic-to-real domain gaps through a multi-branch adaptive architecture. Our approach integrates MFIBA, CSAM, and MFAFM to achieve comprehensive feature learning, precise color restoration, and context-aware haze suppression. The introduction of RRSHID dataset—the first large-scale real-world benchmark—provides critical support for modeling complex atmospheric variations and sensor-specific degradations. Extensive validation demonstrates that MCAF-Net not only outperforms existing methods in real-world scenarios but also maintains competitive efficiency with minimal computational overhead, offering practical value for time-sensitive earth observation tasks. While the current framework excels in static haze removal, challenges remain in handling dynamic atmospheric interactions and multi-modal sensor data. Future efforts will focus on integrating temporal modeling and cross-spectral fusion to advance toward all-weather, physics-aware RS restoration systems. Additionally, we recognize that the ultimate value of RSID lies in enhancing the performance of downstream tasks. The RRSHID dataset and MCAF-Net method have the potential to significantly improve applications such as object detection, change detection, and land cover classification under hazy conditions. Haze typically reduces the accuracy of these tasks, especially for small objects and low-contrast features. Our preliminary analysis indicates that using MCAF-Net as a preprocessing step can improve the performance of these tasks. In future work, we plan to construct benchmark datasets with specific task annotations to quantitatively evaluate the impact of dehazing on downstream applications, further validating the practical value of our proposed method in real-world RS applications.